The weights of trout in a certain river follow a normal distribution with mean 3 pounds and standard deviation 0.8 pounds. Fishermen must throw back any trout under the minimum weight of 2 pounds. What percent will be thrown back?
Answer by stanbon(75887)

The weights of trout in a certain river follow a normal distribution with mean 3 pounds and standard deviation 0.8 pounds. Fishermen must throw back any trout under the minimum weight of 2 pounds. What percent will be thrown back?
-------
z(2) = (2-3)/0.8 = -1/0.8 = -1.25
-------
P(x < 2) = P(z < -1.25) = normalcdf(-100,-1.25) = 0.1056 = 10.56%
==================
cheers,
Stan H.